OpenStudy (kiiraa_x3):

HELP PLEASE TRYING TO GRADUATE !!! HELPPPPPPPPPPPPPPP 1.Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2.Using complete sentences, explain how to convert your birthday function into standard form. 3.Graph your function. Include your graph below. 4.Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (kiiraa_x3):

@Nnesha @vera_ewing @PoetryPrincess @mattyboyy @adamaero @chlobohoe

OpenStudy (kiiraa_x3):

@InfiniteMinds @TrojanPoem

OpenStudy (kiiraa_x3):

@sweetburger @sophiecat4 @ElisaNeedsHelp

OpenStudy (kiiraa_x3):

I already have the first one done I just need number 2.,3 and 4

OpenStudy (kiiraa_x3):

@InfiniteMinds :(

OpenStudy (kiiraa_x3):

@wmj259 @amccall98 @staldk3

OpenStudy (anonymous):

mh. im sorry, my computer lagged

OpenStudy (kiiraa_x3):

its ok can you help ? :(

OpenStudy (kiiraa_x3):

@InfiniteMinds

OpenStudy (kiiraa_x3):

ANYONE PLEASE!!! iM IN DESPERATE NEED OF HELP

OpenStudy (kiiraa_x3):

HELLO

OpenStudy (kiiraa_x3):

@Nurali

OpenStudy (kiiraa_x3):

please help :( @Nurali

OpenStudy (kiiraa_x3):

I looked at that a long time ago , but I don't understand how to do it sine I have a different bithday , so it would be a different equation :( @Nurali

OpenStudy (kiiraa_x3):

@Nurali @Nnesha @Librarian

Nnesha (nnesha):

hm let's get started ! so do u know the quadratic equation in vertex form ?

Nnesha (nnesha):

@kiiraa_x3

OpenStudy (kiiraa_x3):

what do you mean?

Nnesha (nnesha):

general form of quadratic equation ?

OpenStudy (kiiraa_x3):

like for number 1??

Nnesha (nnesha):

vertex form of quadratic equation is \[\huge\rm y = a (x - h)^2 + k\] 1st) given a should be between 4 and -4 so choose 1 a =1 (easy to deal) h = birth month k= birth date replace variables by their values

OpenStudy (kiiraa_x3):

My bday is December 9. so, f(x)=1(x-12)^2+9. Right?

OpenStudy (kiiraa_x3):

now I need help with number 2.

OpenStudy (kiiraa_x3):

@Nnesha

OpenStudy (kiiraa_x3):

hello!!!! @Nurali @Nnesha

Nnesha (nnesha):

yes that's right now (x -12)^2 can be written as (x-12)(x-12) foil it out standard form of quadratic equation is \[\huge\rm y = Ax^2 +Bx +C\]

OpenStudy (kiiraa_x3):

how do I do that

Nnesha (nnesha):

 multiply which is in 1st parentheses by 2nd one